### v5

# ARTalk: Speech-Driven 3D Head Animation via Autoregressive Model

Abstract.
Speech-driven 3D facial animation aims to generate realistic lip movements and facial expressions for 3D head models from arbitrary audio clips. Although existing diffusion-based methods are capable of producing natural motions, their slow generation speed limits their application potential. In this paper, we introduce a novel autoregressive model that achieves real-time generation of highly synchronized lip movements and realistic head poses and eye blinks by learning a mapping from speech to a multi-scale motion codebook. Furthermore, our model can adapt to unseen speaking styles, enabling the creation of 3D talking avatars with unique personal styles beyond the identities seen during training. Extensive evaluations and user studies demonstrate that our method outperforms existing approaches in lip synchronization accuracy and perceived quality.
Speech-driven motion generation, Head animation, Autoregressive models
submissionid: 1975journal: TOGjournalyear: 2025copyright: acmlicensedconference: SIGGRAPH Asia 2025 Conference Papers; December 15–18, 2025; Hong Kong, Hong Kongbooktitle: SIGGRAPH Asia 2025 Conference Papers (SA Conference Papers ’25), December 15–18, 2025, Hong Kong, Hong Kongdoi: 10.1145/3757377.3763955isbn: 979-8-4007-2137-3/2025/12ccs: Computing methodologies Computer graphics
1. Introduction
Figure 1. We present ARTalk, a framework for speech-driven 3D facial motion generation. Our method learns a mapping from speech to a multi-scale motion code, enabling the real-time generation of realistic and diverse animation sequences.
[…]
A concept pipeline of our method.
The rapid evolution of large language models (Achiam et al., 2023; Team et al., 2023; Guo et al., 2025) in recent years has driven significant progress in artificial intelligence, particularly in generating human-like text and facilitating complex conversational interactions. However, many such systems primarily rely on text or synthesized speech for communication, often lacking the visual embodiment necessary for achieving rich and intuitive human-computer interaction. To bridge this gap and create more engaging interactive experiences, research into digital humans has garnered widespread attention from both academia and industry. Due to the intrinsic and natural correlation between the speech and facial expressions, speech-driven 3D facial animation has emerged as a key component for shaping lifelike virtual characters. Therefore, advancements in this field are essential for promoting the application of digital humans in various scenarios such as education and entertainment.
Speech-driven 3D head motion generation aims to synthesize natural and temporally synchronized facial expressions, particularly lip movements, along with realistic head motions directly from speech input. Recent years have witnessed significant advancements in speech-driven motion generation. Autoregressive models, for instance, have been explored to predict deformation offsets on 3D meshes (Richard et al., 2021; Xing et al., 2023; Sung-Bin et al., 2024; Nocentini et al., 2024). However, the inherent complexity of representing facial motion through mesh deformations in 3D space often struggles to fully capture the intricate and non-linear mappings between speech and the resulting nuanced facial expressions. This limitation frequently leads to over-smoothed non-lip regions and a lack of fine-grained realism. More recently, diffusion models (Ho et al., 2020; Tevet et al., 2023; Stan et al., 2023; Sun et al., 2024) have demonstrated compelling generative capabilities, including in the domain of speech-driven motion. For example, FaceDiffuser (Stan et al., 2023) leverages diffusion processes to generate motion directly on 3D meshes, while DiffPoseTalk (Sun et al., 2024) utilizes diffusion on the parametric FLAME model (Li et al., 2017) blendshapes to produce natural facial expressions and head movements. Despite their impressive results in generating high-fidelity motions, these diffusion-based approaches typically suffer from high computational costs, hindering their applicability in real-time and low-latency scenarios. In parallel, multi-scale autoregressive models (Tian et al., 2024) have achieved notable success in image generation tasks, offering a balance between generation speed and output quality. However, these architectures are primarily designed for processing single, multi-scale static samples rather than sequential temporal data, making their direct application to the temporal dynamics inherent in speech-to-motion generation challenging.
To overcome the limitations of existing methods and achieve high-quality, natural, and fast speech-driven motion generation, we propose a novel approach based on temporal windows and multi-scale autoregression. Our method segments the speech input into consecutive temporal windows for motion generation, a window-wise processing strategy crucial for achieving low-latency motion synthesis. Within each temporal window, we further employ a multi-scale encoding and generation strategy to capture facial expressions and head movements with rich details. For instance, for a 4-second temporal window containing 100 motion frames, we encode it into a set of multi-scale representations with varying temporal resolutions, such as encodings with lengths of [1, 5, 25, 50, 100], and adopt a similar multi-scale decoding process during generation. This design allows the model to simultaneously capture global temporal context and fine-grained details of each motion frame, resulting in more expressive motions. To minimize encoding errors and ensure temporal coherence of the generated motions, we extend a VQ codec to encode multi-scale motion codes across two consecutive temporal windows, thereby effectively modeling and transferring temporal dependencies between windows. Additionally, we design a novel autoregressive generator that can autoregressively predict the multi-scale motion codes for the current temporal window conditioned on the speech features of the current window and the previously generated motion sequence. This conditional generation mechanism enables our model to achieve high-quality multi-scale motion generation within each temporal window while maintaining smooth transitions and temporal consistency across windows. Furthermore, we integrate a Transformer-based style encoder to extract style representations from example motion clips. This allows our model to generate stylized motions with unique individual characteristics and helps to disentangle the complex many-to-many mapping between speech and motion. Through these innovations, our method can generate personalized and high-fidelity lip movements, facial expressions, and head motions, demonstrating significant potential for various speech-driven downstream applications.
Our main contributions are as follows:
  • We propose ARTalk, a novel autoregressive framework capable of generating natural 3D facial motions with head poses in real time.
  • We design a novel encoder-decoder that encode motions from consecutive time windows to produce temporally dependent multi-scale motion representations.
  • We introduce a novel conditional autoregressive generator on temporal and feature scales, enabling motion generations tightly aligned with speech conditions and consistent across time windows.
2. Related Work
Table 1. Comparison across methods, including style adaptation during inference (w/ Style), head pose generation (w/ Pose), and real-time capability (Real-time). ARTalk (ours) is the only method to achieve all three features, demonstrating its comprehensive advantages over baseline methods.
Method w/ Style w/ Pose Real-time
FaceFormer (Fan et al., 2022)
CodeTalker (Xing et al., 2023)
SelfTalk (Peng et al., 2023a)
FaceDiffuser (Stan et al., 2023)
MultiTalk (Sung-Bin et al., 2024)
ScanTalk (Nocentini et al., 2024)
UniTalker (Fan et al., 2024)
DiffPoseTalk (Sun et al., 2024)
ARTalk (Ours)
Figure 2. ARTalk involves two separated parts. (a) shows our temporal multi-scale VQ autoencoder. It encodes motion sequences into multi-scale token maps $M_{k_{1}},M_{k_{2}},...,M_{K}]$ using a shared codebook and causal masking on temporal. (b) shows The ARTalk Causal Transformer, where training uses ground truth tokens with a block-wise causal attention mask, and inference autoregressively predicts motion tokens conditioned on speech features and last scale tokens and last time window motions.
[…]
Method pipeline of our method.
2.1. Speech-Driven 3D Facial Animation
Research on audio-driven 3D motion generation has been an area of interest for decades, with methods evolving significantly over time. Early approaches (Taylor et al., 2012; Xu et al., 2013; Edwards et al., 2016) relied primarily on procedural methods, which segmented speech into phonemes and mapped them to predefined visemes through handcrafted rules. Although procedural methods provide explicit control over the generated animations, they typically require complex parameter tuning and fail to capture the diversity and complexity of real-world speaking styles. As a result, procedural methods struggled to deliver animations that appeared both natural and adaptable to varying speech dynamics. In recent years, learning-based approaches (Fan et al., 2022; Xing et al., 2023; Sun et al., 2024; Daněček et al., 2023; Peng et al., 2023b; Yang et al., 2024; Fan et al., 2024; Chae-Yeon et al., 2025) have advanced rapidly, addressing these limitations and enabling more natural and expressive facial animations. Additionally, some approaches (Yi et al., 2022; Ye et al., 2024; Zhang et al., 2023; Tan et al., 2024) focus on directly generating talking head videos without explicitly modeling motion. However, this limits their ability to integrate with motion-driven downstream applications, restricting their broader applicability.
2.2. Autoregressive 3D Facial Motion Generation
Autoregressive (AR) methods model the temporal sequence of facial motion in a step-by-step manner. These methods typically leverage pre-trained speech model features (Baevski et al., 2020; Hsu et al., 2021; Défossez et al., 2024) as speech representations, which are subsequently mapped to 3D deformable model parameters or 3D meshes via neural networks. For example, FaceFormer (Fan et al., 2022) autoregressively predicts the continuous facial motion parameters while MeshTalk (Richard et al., 2021) learns a prior over categorical expression tokens. CodeTalker (Xing et al., 2023) and MultiTalk (Sung-Bin et al., 2024) tokenize facial motions with a VQ-VAE and train a transformer decoder to predict subsequent motion tokens based on audio features. Learn2Talk (Zhuang et al., 2024) augments FaceFormer with lip-sync and lip-reading losses, improving alignment with speech. MMHead (Wu et al., 2024) provides text descriptions as an extra condition for controlling the generated motion. However, these methods also show limitations. Temporal autoregressive modeling often under represents motion within each time window, leading to overly smooth lip movements and failing to capture the complex speech-to-motion mapping. This issue becomes more pronounced with larger and more diverse datasets. Additionally, many approaches (Fan et al., 2022; Xing et al., 2023; Sung-Bin et al., 2024) rely on a predefined set of style labels to reduce the complexity of the mapping, which limits their ability to adapt to new individuals and styles.
2.3. Diffusion based 3D Facial Motion Generation
Diffusion models (Ho et al., 2020; Zhu et al., 2023; Alexanderson et al., 2023; Stan et al., 2023; Aneja et al., 2024; Baevski et al., 2020; Ma et al., 2024; Sun et al., 2024; Zhang et al., 2024) have recently gained traction for generative tasks because of their strong modeling capacities. FaceDiffuser (Stan et al., 2023) conditions a diffusion model on audio features to predict displacements from a neutral template. Scantalk (Nocentini et al., 2024) uses a DiffusionNet (Sharp et al., 2022) structure to overcome the fixed topology limitation and perform diffusion on arbitrary meshes. Similarly, DiffPoseTalk (Sun et al., 2024) uses a diffusion-based transformer decoder to generate both expression and pose (blendshapes) conditioned on audio and learned style features. Although these methods generate high-fidelity and realistic facial motion, their iterative sampling steps can be computationally expensive and may limit real-time applicability.
In this paper, we leverage FLAME (Li et al., 2017) as our motion representation and propose a novel autoregressive framework for speech-driven motion generation. It not only outperform current diffusion models in generating facial motion, but also achieves real-time speed and produces natural head poses. A comparison of our method with existing approaches is shown in Table 1. Our method enables arbitrary style inference, allowing stylized motion generation at inference time by providing an example motion clip.
3. Method
We provide an overview of our method in Figure 2. We adopt the widely used 3DMM (Li et al., 2017) as the facial representation and leverage a multi-scale VQ autoencoder model to train and obtain a multi-scale motion codebook along with its corresponding encoder-decoder, enabling a discrete representation of the motion space. Subsequently, we train a multi-scale autoregressive model to map speech information to the discrete motion space.
In the following subsections, Section 3.1 details the problem definition and the 3DMM representation, Section 3.2 explains the multi-scale motion VQ autoencoder model, Section 3.3 introduces the multi-scale autoregressive model.
3.1. Preliminaries
We adopt the widely used 3D morphable model (3DMM) FLAME (Li et al., 2017) to represent facial motion, where each frame is parameterized by shape $\beta$, expression $\psi$, and pose $\theta$. Given these parameters, the 3D face mesh, consisting of 5,023 vertices, is reconstructed through blend shape and rotation operations. Compared to previous methods (Fan et al., 2022; Xing et al., 2023; Peng et al., 2023a; Nocentini et al., 2024) that directly model the mesh, FLAME simplifies motion modeling by converting complex vertex movements into lower-dimensional blend shape parameters while preserving expression and motion details.
We define the motion vector $\mathbf{M}\in\mathbb{R}^{K\times D}$ as the concatenation of per-frame expression and pose parameters over $K$ frames. The corresponding vertex coordinates $\mathbf{V}\in\mathbb{R}^{K\times N\times 3}$ store the 3D positions of $N$ mesh vertices per frame, where $\mathbf{V}_{\mathrm{lips}}\subset\mathbf{V}$ denotes the subset corresponding to the lip region, which is particularly important for speech related modeling. Further details on 3DMM can be found in the supplementary material.
3.2. Temporal Multi-scale VQ Autoencoder
Predicting motion frames from speech is a discrete task involving highly dense temporal sequences and complex mappings. Inspired by the success of discrete representations in image and motion generation (Xing et al., 2023; van den Oord et al., 2017; Zhou et al., 2022), the effectiveness of multiscale audio modeling for animation generation (Jung et al., 2024) and the effectiveness of multi-scale encoding in (Tian et al., 2024), we propose a temporal multi-scale VQ autoencoder to efficiently model motion dynamics.
The input motion sequence, consisting of $K$ frames, is first processed by a Transformer encoder, which extracts temporal features and maps them into a latent space. These representations are then quantized using a multi-scale codebook that captures motion information at varying temporal resolutions $\bigl{[}k_{1},k_{2},\ldots,K\bigr{]}$, producing discrete tokens. The Transformer decoder reconstructs the motion from these discrete embeddings, attending to both local and long-range dependencies.
Multi-scale residual VQ. To progressively refine motion representations, we adopt a residual vector quantization approach:
(1) $\mathbf{h}^{(l)}$ $=\mathrm{Interp}\bigl{(}\mathrm{Quant}(\mathbf{r}^{(l-1)}),\,k_{l}\bigr{)},$
$\mathbf{r}^{(l)}$ $=\mathbf{r}^{(l-1)}-\mathbf{h}^{(l)},$
for $l=1,\ldots,L$, where $\mathbf{r}^{(0)}$ is the encoder output. The function $\mathrm{Quant}(\cdot)$ assigns features to the closest codebook entries, while $\mathrm{Interp}(\cdot,k_{l})$ adjusts resolutions to ensure smooth transitions. By accumulating the outputs $\mathbf{h}^{(l)}$ across scales, the model retains high fidelity while removing redundancy in a structured manner.
Residual accumulation at the finest scale $K$ mitigates information loss, while downsampling via region-based interpolation enhances efficiency. We use a shared codebook across scales to maintain a unified motion space and simplify autoregressive modeling. This strategy improves temporal consistency and preserves stylistic coherence across different resolutions.
Temporal causal reasoning. To ensure stability over extended sequences, we incorporate a causal reasoning mechanism that enforces cross-window consistency. Our approach considers two consecutive time windows, T1T-1 and $T$. During training and inference, causal masks are applied, ensuring that predictions for $T$ depend only on previously observed information from T1T-1, without leaking future details. This prevents temporal discontinuities and enhances the smoothness of long-term motion.
By integrating multi-scale residual quantization with causal reasoning and Transformer-based encoding/decoding, our method ensures both fine-grained expressiveness and long-term coherence in generated motion sequences.
Training objectives.
To train the VQ autoencoder, we use a hybrid loss function balancing motion accuracy, temporal smoothness, and codebook stability.
The reconstruction loss enforces alignment between predicted motion vectors $M^{\prime}$ and the ground truth $M$, with additional constraints on lip and facial vertices:
(2) $L_{\text{recon}}=\|\hat{M}-M\|_{1}+w_{\text{lips}}\|\hat{V}_{\text{lips}}-V_{\text{lips}}\|^{2}+\|\hat{V}-V\|^{2}$
To ensure smooth transitions, we penalize velocity and acceleration differences:
(3) $L_{\text{vel}}$ $=\|(\hat{V}_{1:}-\hat{V}_{:-1})-(V_{1:}-V_{:-1})\|^{2},$
(4) $L_{\text{smooth}}$ $=\|\hat{V}_{2:}-2\hat{V}_{1:-1}+\hat{V}_{:-2}\|^{2}.$
We also apply the standard VQ losses which we term as $L_{\text{cb}}$ to encourage stable VQ assignments and prevent mode collapse.
The final training objective is:
(5) $L_{\text{VQ}}=L_{\text{recon}}+\lambda_{\text{vel}}L_{\text{vel}}+\lambda_{\text{smooth}}L_{\text{smooth}}+L_{\text{cb}},$
where $\lambda_{\text{vel}}$ and $\lambda_{\text{smooth}}$ control the temporal constraints.
By jointly optimizing these terms, the model learns compact yet expressive motion representations, ensuring high-fidelity synthesis with strong temporal coherence.
3.3. Speech-to-Motion Autoregressive Model
After training the VQ autoencoder (Section 3.2), we obtain discrete codes $\{\mathbf{z}_{T}^{(l)}\}$ at multiple resolutions $k_{l}$ for each time window $T$, where $l=1,\dots,L$ spans from coarsest to finest. To ensure long-term temporal coherence, we model these codes with a Transformer that is autoregressive across scales $(1\!\to\!L)$ and across adjacent windows $(T-1\!\to\!T)$.
Architecture. Figure 2(b) illustrates the AR model, where orange arrows mark training-only components. A pre-trained HuBERT (Hsu et al., 2021) encoder extracts speech features $\mathbf{a}_{T}$, resampled to each scale $k_{l}$. A style token $\mathbf{s}$ encodes speaker identity (Sun et al., 2024). Given the previous window’s finest-scale codes $\mathbf{z}_{T-1}^{(L)}$, we predict $\mathbf{z}_{T}^{(1)},\dots,\mathbf{z}_{T}^{(L)}$ for the current window.
Two-level autoregression. Within each window $T$, the codes are generated scale by scale:
(6) $p(\{\mathbf{z}_{T}^{(l)}\}_{l=1}^{L}\mid\mathbf{z}_{T-1}^{(L)},\mathbf{a}_{T},\mathbf{s})\;=\;\prod_{l=1}^{L}p\!\bigl{(}\mathbf{z}_{T}^{(l)}\mid\mathbf{z}_{T}^{(<l)},\mathbf{z}_{T-1}^{(L)},\mathbf{a}_{T},\mathbf{s}\bigr{)},$
where $\mathbf{z}_{T}^{(<l)}$ denotes codes at all coarser scales. Inside each scale $l$, tokens are predicted in parallel (blockwise) while attending causally to previously generated scales and the previous window.
Training and inference. During training, ground-truth codes $\{\mathbf{z}_{T}^{(l)\!\text{(gt)}}\}$ supervise the model via cross-entropy:
(7) $L_{\mathrm{AR}}$ $=\;-\sum_{T}\sum_{l=1}^{L}\log\,p(\mathbf{z}_{T}^{(l)\!\text{(gt)}}\mid\mathbf{z}_{T}^{(<l)\!\text{(gt)}},\mathbf{z}_{T-1}^{(L)\!\text{(gt)}},\mathbf{a}_{T},\mathbf{s}).$
At inference, we sample or select each scale’s codes in sequence, conditioned on $\mathbf{z}_{T-1}^{(L)}$, $\mathbf{a}_{T}$, and $\mathbf{s}$. Repeating this process over windows yields motion sequences that remain temporally coherent and stylistically consistent throughout.
Figure 3. Comparison of efficiency and performance across different methods.
4. Experiments
Datasets. We use the TFHP dataset proposed by  Sun et al. (2024) to train our model, which consists of 1,052 video clips from 588 subjects, with a total duration of approximately 26.5 hours. All videos are tracked at 25 fps, resulting in approximately 2,385,000 action frames. For training and evaluation, we adopt the train/test split provided in the original paper (Sun et al., 2024). Furthermore, we evaluate the generalization performance of our model on the test split of the widely used VOCASET dataset (Cudeiro et al., 2019). The test split consists of 80 audiovisual sequence pairs from 2 subjects. To adapt VOCASET to our method, we track the FLAME parameters at 25 fps and conduct evaluation on our tracked data. It is worth noting that we don’t use VOCASET to train or fine-tune our method.
Implementation Details. Our method is implemented on the PyTorch framework (Paszke et al., 2017). In the first stage, we train our VQ autoencoder to obtain a multi-scale motion codebook. The motion codebook consists of 256 entries, each with a code dimension of 64. The time window size is 100 frames (4 seconds), and the multi-scale levels are [1, 5, 25, 50, 100]. At this stage, we used the AdamW optimizer with a learning rate of 1.0e-4 and a total batch size of 64 for 50,000 iterations. In the second stage, we train the multi-scale autoregressive model using the AdamW optimizer with the same learning rate of 1.0e-4 and a batch size of 64 for 50,000 iterations. During this stage, we employ a frozen pre-trained HuBERT (Hsu et al., 2021) backbone.
All training was conducted on one NVIDIA Tesla A100 GPU, requiring a total of approximately 13 GPU hours (5 hours for the first stage and 8 hours for the second stage), demonstrating efficient training resource utilization. During inference, our method needs only 0.01 seconds to generate motions for 1 second on an NVIDIA Tesla A100 GPU and 0.057 seconds on an Apple M2 Pro chip, showcasing high inference efficiency and low latency. For more implementation details, please refer to the supplementary materials.
Table 2. Quantitative results on the TFHP (Sun et al., 2024) dataset. We use colors to denote the first and second places respectively. {*} indicates that the method was not trained on TFHP, while all other methods were trained or fine-tuned on TFHP.
Method LVE $\downarrow$ FFD $\downarrow$ MOD $\downarrow$
FaceFormer (Fan et al., 2022) 12.72 22.06 2.73
CodeTalker (Xing et al., 2023) 11.78 20.39 2.43
SelfTalk (Peng et al., 2023a) 12.07 23.74 2.57
FaceDiffuser (Stan et al., 2023) 11.92 22.17 2.55
MultiTalk (Sung-Bin et al., 2024) 12.23 24.42 2.48
ScanTalk (Nocentini et al., 2024) 12.14 21.02 3.20
UniTalker (Nocentini et al., 2024) 11.61 29.31 2.07
DiffPoseTalk (Sun et al., 2024) 10.39 20.15 2.07
ARTalk (Ours) 9.34 18.15 1.81
Table 3. Quantitative results on the VOCASET-Test (Cudeiro et al., 2019) dataset. We use colors to denote the first and second places respectively. {*} indicates that the method was not trained on VOCASET. It is worth noting that our method is not trained or fine-tuned on VOCASET.
Method LVE $\downarrow$ FFD $\downarrow$ MOD $\downarrow$
FaceFormer (Fan et al., 2022) 8.28 15.62 1.89
CodeTalker (Xing et al., 2023) 7.73 18.86 1.81
SelfTalk (Peng et al., 2023a) 7.71 28.70 1.79
FaceDiffuser (Stan et al., 2023) 8.00 21.46 1.84
MultiTalk (Sung-Bin et al., 2024) 12.33 25.11 2.82
ScanTalk (Nocentini et al., 2024) 7.15 15.56 1.61
UniTalker (Nocentini et al., 2024) 7.38 21.72 1.89
DiffPoseTalk (Sun et al., 2024) 10.01 20.64 2.29
ARTalk (Ours) 7.57 15.49 1.78
Figure 4. Qualitative comparison with existing methods (all head poses fixed). The first four rows are from the TFHP dataset, and the last two rows are from the VOCASET dataset. Our method shows better alignment with the ground truth in expression style, mouth dynamics, and lip synchronization. Additional results are available in the supplementary materials and demo videos.
[…]
Qualitative comparison.
4.1. Quantitative Results
Based on previous studies (Richard et al., 2021; Fan et al., 2022; Xing et al., 2023; Nocentini et al., 2024; Sun et al., 2024), we adopted two quantitative metrics, the lip vertex error (LVE) (Richard et al., 2021) and the upper face dynamic deviation (FDD) (Xing et al., 2023), to evaluate the generated facial motions. LVE calculates the maximum L2 error of all lip vertices for each frame, evaluating the largest deviation between predicted and ground truth lip positions. FDD calculates the standard deviation of the motion of each upper facial vertex over time between predictions and ground truth, evaluating the consistency of upper facial motion, which is closely related to speaking styles. Additionally, we use a similar metric as  Sun et al. (2024), mouth opening distance (MOD), to more accurately assess the stylistic similarity of mouth opening movements. MOD measures the average difference in mouth opening region between predictions and ground truth. Compared to LVE, MOD focuses more on the similarity of mouth opening styles and is less sensitive to temporal lip synchronization. For the partitioning of the lip area and upper face, we used the mask partitioning officially provided by FLAME (Li et al., 2017). The lip region contains 254 points, while the upper face, including eye region and forehead, contains 884 points.
We present the quantitative comparison on TFHP dataset (Sun et al., 2024) in Table 2. The baseline methods FaceFormer (Fan et al., 2022), CodeTalker (Xing et al., 2023), SelfTalk (Peng et al., 2023a), and FaceDiffuser (Stan et al., 2023) are mesh-based approaches and cannot generalize to arbitrary styles or new meshes. To ensure a fair comparison, we retrained these methods on meshes generated from the TFHP dataset. For MultiTalk (Sung-Bin et al., 2024), which is trained on FLAME meshes and supports language-based stylization, we used its English style for evaluation. ScanTalk (Nocentini et al., 2024) claims to work on arbitrary meshes, so we use the officially provided pre-trained weights. For UniTalker (Fan et al., 2024), we compute the metrics using the meshes generated by the generalized pivot identities.
For DiffPoseTalk (Sun et al., 2024), we used pre-trained weights on the TFHP dataset (including head pose) for evaluation. The results in the Table 2 show that our method achieving significant improvements in lip synchronization accuracy (LVE) and style alignment (FFD and MOD), indicating that our method not only achieves precise lip synchronization but also effectively captures personalized speaking styles.
To demonstrate the generalizability of our method, we evaluated it on the VOCASET dataset (Cudeiro et al., 2019). The results are presented in Table 3. Notably, for some baseline methods (Fan et al., 2022; Xing et al., 2023; Peng et al., 2023a; Stan et al., 2023; Nocentini et al., 2024), we used their pre-trained weights on VOCASET, whereas our method was not trained or fine-tuned on this dataset. Despite this, our method achieved highly competitive performance and outperformed most baseline methods specifically trained on this dataset. For MultiTalk (Sung-Bin et al., 2024) and DiffPoseTalk (Sun et al., 2024) which designed with style generalization capabilities, we directly testing it on the VOCASET dataset without fine-tuning. The results indicate that our method surpasses them in terms of generalization ability, further verifying the robustness of our method in handling unseen styles and data.
We also present a comparison of LVE and efficiency in Figure 3. Although our model adopts a two-level autoregressive framework, it remains more efficient than fully autoregressive methods. This efficiency is achieved by utilizing longer window lengths while retaining only the past 4 seconds of motion frames. Additionally, frames within each scale are generated in parallel within the window, further improving computational efficiency.
4.2. Qualitative Results
In Figure 4, we present a qualitative comparison between our method and other baseline approaches. Our method demonstrates excellent lip synchronization, accurately capturing various phonetic elements. Furthermore, the generated results exhibit realistic facial expressions and mouth opening, closely matching the style of the ground truth. Notably, our method also generates realistic blinking and head movements, which are implicitly learned and encoded within the motion codebook. In Figure 5, we also show the ability of our method on head movements. It can capture accents well and generate reasonable head movements. Additional qualitative evaluation results are available in the supplementary videos.
Table 4. User study results. The percentages represent the proportion of users who preferred the our method over the baseline in each category. The values in brackets represent the 95% confidence interval, indicating that we are 95% confident that the true value of the performance metric lies within this range. Among them, Sync represents lip synchronization, Style represents the consistency of style, N-Exp and N-Pose represent the facial expression naturalness and head pose naturalness respectively.
Method Sync (%) N-Exp (%) Style (%) N-Pose (%)
FaceFormer 75.0 [70.4, 79.6] 89.3 [86.3, 92.8] 88.1 [85.0, 91.8] -
CodeTalker 84.5 [80.7, 88.4] 86.9 [83.3, 90.5] 92.9 [90.5, 95.9] -
SelfTalk 85.7 [82.0, 89.5] 89.3 [86.3, 92.8] 91.7 [89.1, 94.9] -
FaceDiffuser 88.1 [85.0, 91.8] 90.4 [87.3, 93.6] 86.9 [83.3, 90.5] -
MultiTalk 78.6 [74.5, 83.2] 76.2 [72.0, 81.0] 78.6 [74.5, 83.2] -
ScanTalk 88.1 [85.0, 91.8] 90.5 [87.7, 93.9] 90.5 [87.7, 93.9] -
UniTalker 81.8 [77.7, 86.0] 83.6 [79.7, 87.6] 81.3 [77.4, 85.7] -
DiffPoseTalk 63.1 [58.2, 68.5] 59.5 [54.3, 64.8] 60.7 [55.5, 65.9] 58.3 [53.1, 63.6]
4.3. User Study
User studies are a reliable approach for evaluating 3D talking heads. To comprehensively compare our method with baseline methods, we conducted a user study focusing on four key metrics: lip synchronization, facial expression naturalness, style consistency and head pose naturalness.
Given that head movements can significantly influence users’ evaluation of lip synchronization, the study was divided into two parts. In the first part, videos with fixed head poses were rendered across all methods to assess lip synchronization and facial expression realism and style consistency. In the second part, videos with dynamic head poses were rendered to evaluate style consistency and head pose naturalness. All comparisons were conducted using pairwise comparisons, where motions generated by our method and a competing baseline were displayed side by side, along with the ground truth provided as a reference for users. After watching the videos, users selected the animation they perceived as better based on their subjective preference. The proportion of user selections was calculated to quantify satisfaction.
As presented in Table 4, our method significantly outperformed the baseline methods in lip synchronization, style consistency and facial expression realism. Furthermore, our method achieved superior head pose naturalness compared to DiffPoseTalk (Sun et al., 2024), demonstrating its comprehensive advantages.
4.4. Ablation study
Multi-Scale Autoregression. To validate the importance of multi-scale autoregressive modeling within a single time window, we conducted experiments by removing the multi-scale autoregressive mechanism entirely. The results shown in Table 5 indicate a significant drop in generation accuracy without multi-scale autoregression. We attribute this decline to the model’s inability to fully capture the complex details of the speech-to-motion mapping without the hierarchical structure provided by multi-scale modeling.
Figure 5. Qualitative results of head pose. When certain words are stressed or when accents occur, the model produces nodding motions similar to human behavior.
Temporal Encoder and Temporal Autoregression. To assess the necessity of temporal encoding and autoregression across time windows, we replaced our proposed temporal encoder and autoregressive model with a standard multi-scale encoder and a single-window multi-scale autoregressive model, respectively. As shown in Table 5, the absence of temporal modeling in VQ autoencoder or autoregressive both result in degraded generation quality. Additionally, we also observed temporal discontinuities in the motion sequences, further emphasizing the critical role of our temporal design in achieving smooth and coherent outputs.
Speaker Style Embedding. We also evaluated the effect of removing the speaker style feature, which serves as the starting condition for the autoregressive process. Without this feature, the complexity of the many-to-many mapping between speech and motion increases significantly, and the generated motions lack personalized style. As reflected in Table 5, the removal of style features causes a substantial decline in generation quality, highlighting their importance in achieving stylistic and expressive outputs.
Key Hyperparameter Choices. We explored different number of motion frames within a single time window, which is critical for our method. Specifically, we tested window lengths of 8, 25, 50 frames. For a window length of 8, we used a multi-scale sequence of [1,2,4,8], while for 25 and 50, we adopted [1,5,10,15,25] and [1,5,10,25,50], respectively. The results, presented in Table 5, demonstrate that the choice of window size significantly impacts performance. However, the window size depends heavily on the downstream task characteristics, such as whether the input consists of streaming audio chunks. In general, longer temporal windows generally produce better results, but longer windows mean that more audio needs to be collected to output motion. In this paper, we chose a window length of 100 frames and a multi-scale sequence of [1, 5, 25, 50, 100] as it is consistent with previous work such as DiffPoseTalk (Sun et al., 2024) and achieves a good balance between quality and efficiency. Further exploration of the number of multi-scale layers and audio chunk lengths is provided in the supplementary material.
Integration with Downstream Tasks Our method can be integrated into various FLAME-based downstream applications, enabling a wider range of use cases. While this extends beyond the primary scope of this paper, we also demonstrate its application in GAGAvatar (Chu and Harada, 2024). As shown in Figure 6, our method can generate control signals for GAGAvatar, effectively transforming it into a speech-driven one-shot dynamic avatar reconstruction method.
Table 5. Ablation results on TFHP dataset.
[…]
w/o Multi-Scale AR 14.14 22.04 3.06
w/o Temporal VQ 9.82 18.60 1.86
w/o Temporal AR 9.97 18.71 1.99
w/o Style Embedding 11.80 21.46 2.37
Clip Length 8 Layer 5 11.73 24.89 2.25
Clip Length 25 Layer 5 10.20 19.00 1.97
Clip Length 50 Layer 5 9.78 18.03 1.89
ARTalk (Full) 9.34 18.15 1.81
Figure 6. We reconstruct the avatar using GAGAvatar (Chu and Harada, 2024) and drive it with our ARTalk model, enabling speech-driven dynamic avatar generation.
5. Discussion and Conclusion
In this paper, we introduced ARTalk, a novel framework for generating 3D facial and head motions from speech. The core innovation of our method lies in the temporal multi-scale autoencoder and the ARTalk autoregressive Transformer, which together ensure temporal consistency and precise motion generation. Our experimental results demonstrate that ARTalk outperforms state-of-the-art baselines in terms of lip synchronization, expression naturalness and style consistency, while maintaining real-time generation capabilities. We believe that the strong generalization ability of ARTalk make it a promising solution for a wide range of applications, including virtual avatars, language training, and animation production for gaming and movies.
Limitations and future work. While ARTalk demonstrates strong performance in lip synchronization and expressions, a key limitation is that head movements are driven by speech prosody, not semantic context. This prevents the model from generating culturally-specific or meaningful gestures like nodding or shaking the head to indicate affirmation. Additionally, our use of a single style embedding can be challenging for speech with dramatic emotional shifts, which is also a common problem with existing methods. For example, when a speaker suddenly shifts from calm to excited, a single embedding may struggle to accurately capture this transition.
We hope that our work lays a solid foundation for the further research of 3D talking head generation. Future research will focus on addressing the above challenges, especially exploring how to effectively incorporate semantic information into head motion generation and developing more flexible style control methods.
Impact Statement
Given that our method can generate highly realistic talking head sequences, there is a potential risk of misuse, such as deepfake creation and manipulation. To address these concerns, we strongly advocate for watermarking or clearly annotating facial sequences generated by our method as synthetic data. Moreover, we will collaborate on deepfake detection research to further ensure that our proposed method is used for positive and constructive applications. We also aim to raise awareness about these risks and encourage collaboration among governments, developers, researchers, and users to jointly prevent the misuse of synthetic video technologies.
Acknowledgements.
This work was partially supported by JST Moonshot R&D Grant Number JPMJPS2011, CREST Grant Number JPMJCR2015 and Basic Research Grant (Super AI) of Institute for AI and Beyond of the University of Tokyo. In addition, this work was also partially supported by JST SPRING, Grant Number JPMJSP2108.